OpenStudy (anonymous):

In a triangle, the sum of the interior angles is 180o. Find the measure of angle x, if angle y = 43o and angle z = 67o. A. Write the algebraic equation using the information above. (5 points) B. Solve the equation. Show all work. (10 points) C. What is the measure of angle x? Make sure to label your answer. (5 points)

OpenStudy (anonymous):

Is this answer correct? A. Angle y + Angle z + Angle x = 180o Angle y = 43o Angle z = 67o Angle x = X B. 43 + 67 + x = 180 110 + x = 180 110 - 110 + x = 180 - 110 x = 70 43 + 67 + 70 = 180 C. The measure of Angle x is 70 degrees.

OpenStudy (lilg132):

yes it is correct

OpenStudy (anonymous):

Thank you :)